Big Oil, Inc. has a preferred stock outstanding that pays a $5 annual dividend. If investors’ required rate of return is 5 percent, what is the market value of the shares? Round your answer to the nearest cent.

$ _____

If the required return declines to 2 percent, what is the change in the price of the stock? Round your answer to the nearest cent.

The price____ (increases or decreases) by $  _____
Answer #1
Price = Dividend/(interest rate)
= 5/(0.05-0)
= 100

New price

Price = Dividend/(interest rate)
= 5/(0.02-0)
= 250

Price increases by 250-100 = 150